# Resource theories with completely passive states and passive states as free states

### v3

@@ IV Passive states as free states
for all $P\in\Pi$, where $\sigma_{i}=K_{i}\rho K_{i}^{\dagger}/q_{i}$, $q_{i}=\text{Tr}\left\{K^{\dagger}_{i}K_{i}\rho\right\}$, $P_{i}=K_{i}PK_{i}^{\dagger}/\text{Tr}\left\{K^{\dagger}_{i}K_{i}P\right\}$ and $\Lambda(\rho)=\sum_{i}K_{i}\rho K^{\dagger}_{i}=\sum_{i}q_{i}\sigma_{i}$. For instance, Eq. (40) is satisfied if $T(\gamma_{i}|\sigma_{i})\leq T(\gamma|\rho)$ for all $i$, since in general $S(\rho||P)\geq\sum_{i}q_{i}S(\sigma_{i}||P_{i})$. Furthermore, if we consider two systems with Hamiltonian $H_{1}$ and $H_{2}$, and the composite system with Hamiltonian $H=H_{1}+H_{2}$, the ergotropy is super-additive under tensor products [15], i.e., $\mathcal{E}(\rho_{1}\otimes\rho_{2})\geq\mathcal{E}(\rho_{1})+\mathcal{E}(\rho_{2})$, thus it cannot be considered as a measure.
For $d=2$, all the passive states are completely passive states, i.e., $P=\gamma_{\beta(P)}$ with $\beta(P)=1/T(P)$ for all $P\in\Pi$, and thus $\Pi=C$. We note that $T(P|\rho)=T(P)$ does not depend on $\rho$, so that from Eq. (30) we get $\mathcal{E}(\rho)=F(\rho)$. However, for $d>2$, there are free operations $\Lambda$ that do not increase the temperature for a fixed selection, i.e., $\Lambda\in\mathcal{O}_{p}$ such that $T(\Lambda(P)|\rho)\leq T(P|\rho)$, giving $T(\Lambda(P)|\Lambda(\rho))>T(P|\rho)$. For instance, the dephasing map $\Delta(\rho)$ preserves the passive states, so that satisfies (E,i), and does not increase the ergotropy, since the coherent contribution is nonnegative [16], i.e., $\mathcal{E}_{c}(\rho)=\mathcal{E}(\rho)-\mathcal{E}(\Delta(\rho))\geq 0$. Then, the dephasing map is a free operation, $\Delta\in\mathcal{O}_{p}$. For the dephasing map, we get $T(\Delta(P)|\rho)=T(P|\rho)$, however we can get $T(\Delta(P)|\Delta(\rho))>T(P|\rho)$ for $d>2$. It is enough to consider $d=3$. From an explicit calculation, for $\rho=\frac{1}{2}(|\epsilon_{2}\rangle\langle\epsilon_{2}|+|\epsilon_{3}\rangle\langle\epsilon_{3}|+|\epsilon_{2}\rangle\langle\epsilon_{3}|+|\epsilon_{3}\rangle\langle\epsilon_{2}|)$, we find $P_{\rho}=|\epsilon_{1}\rangle\langle\epsilon_{1}|$, $\Delta(\rho)=\frac{1}{2}(|\epsilon_{2}\rangle\langle\epsilon_{2}|+|\epsilon_{3}\rangle\langle\epsilon_{3}|)$ and $P_{\Delta(\rho)}=\frac{1}{2}(|\epsilon_{1}\rangle\langle\epsilon_{1}|+|\epsilon_{2}\rangle\langle\epsilon_{2}|)$, thus we find that $T(\Delta(P)|\Delta(\rho))>T(P|\rho)$ for the passive states $P=\sum_{k}p_{k}|\epsilon_{k}\rangle\langle\epsilon_{k}|$ with populations $p_{k}\geq p_{k+1}$ satisfying the inequality
@@
We note that if $\Lambda$ is a unital map, then $E(P_{\Lambda(\rho)})\geq E(P_{\rho})$ (see, e.g., Ref. [15]). Thus, if $E(\Lambda(\rho))\leq E(\rho)$, we get
@@
In particular, if $\Lambda(C)\subseteq C$, the condition $E(\Lambda(\gamma_{\beta}))\leq E(\gamma_{\beta})$ is equivalent to $T(\Lambda(\gamma_{\beta}))\leq T(\gamma_{\beta})$, i.e., $\eta_{\Lambda}(C)=1$. If $\Lambda$ is a unitary map defined by a unitary operator $U$, so that $\Lambda(\rho)=U\rho U^{\dagger}$, the condition (E,ii) is equivalent to $E(\Lambda(\rho))\leq E(\rho)$. For instance the map $\Lambda(\rho)=U_{\rho}\rho U_{\rho}^{\dagger}$ that extracts the ergotropy is free. Then, there is no a work locking phenomenon [17, 18, 19], since all the ergotropy can be extracted with free unitary maps.
V Conclusions
@@
Work extraction in thermally isolated quantum systems can be performed through unitary cycles. The maximum work extractable is the ergotropy, which is upper bounded by the free energy. When we have many copies of the system available, the density of maximum work extractable tends to be the free energy. Thus, we defined resource theories where the free energy and the ergotropy are monotones. Completely passive states and passive states are the free states of the resource theories. In particular, we discussed how, similarly to the free energy, the ergotropy is a distance-based monotone. Thus, with the aim to give a unified picture, we expressed both the free energy and the ergotropy in terms of a unique metric, with a non-contractive distance. This distance is defined in terms of the quantum relative entropy and a nonequilibrium temperature, which plays a certain role to characterize the free operations. In detail, any map for which the distance is contractive is a free operation for the corresponding theory. However, there are also free operations for which the distance is not contractive, finding that all the free operations can increase the temperature of the free states at most by a certain factor not smaller than 1.
@@
In conclusion, we hope that our results can be useful to understand how the work extractable can be viewed as a resource and can find applications for the manipulation of the states of work storing centers, e.g., quantum batteries. Furthermore, we believe that our findings can also play some role in understanding what is the temperature of a nonequilibrium state. In particular, the definition of the nonequilibrium temperature is a challenging open problem of quantum thermodynamics, e.g., a nonequilibrium temperature was recently introduced in Ref. [20], which is different from the one introduced by us.

### v4

@@ IV Passive states as free states
for all $P\in\Pi$, where $\sigma_{i}=K_{i}\rho K_{i}^{\dagger}/q_{i}$, $q_{i}=\text{Tr}\left\{K^{\dagger}_{i}K_{i}\rho\right\}$, $P_{i}=K_{i}PK_{i}^{\dagger}/\text{Tr}\left\{K^{\dagger}_{i}K_{i}P\right\}$ and $\Lambda(\rho)=\sum_{i}K_{i}\rho K^{\dagger}_{i}=\sum_{i}q_{i}\sigma_{i}$. For instance, Eq. (40) is satisfied if $T(P_{i}|\sigma_{i})\leq T(P|\rho)$ for all $i$, since in general $S(\rho||P)\geq\sum_{i}q_{i}S(\sigma_{i}||P_{i})$. Furthermore, if we consider two systems with Hamiltonian $H_{1}$ and $H_{2}$, and the composite system with Hamiltonian $H=H_{1}+H_{2}$, the ergotropy is super-additive under tensor products Francica22 , i.e., $\mathcal{E}(\rho_{1}\otimes\rho_{2})\geq\mathcal{E}(\rho_{1})+\mathcal{E}(\rho_{2})$, thus it cannot be considered as a measure.
For $d=2$, all the passive states are completely passive states, i.e., $P=\gamma_{\beta(P)}$ with $\beta(P)=1/T(P)$ for all $P\in\Pi$, and thus $\Pi=C$. We note that $T(P|\rho)=T(P)$ does not depend on $\rho$, so that from Eq. (30) we get $\mathcal{E}(\rho)=F(\rho)$. However, for $d>2$, there are free operations $\Lambda$ that do not increase the temperature for a fixed selection, i.e., $\Lambda\in\mathcal{O}_{p}$ such that $T(\Lambda(P)|\rho)\leq T(P|\rho)$, giving $T(\Lambda(P)|\Lambda(\rho))>T(P|\rho)$. For instance, the dephasing map $\Delta(\rho)$ preserves the passive states, so that satisfies (E,i), and does not increase the ergotropy, since the coherent contribution is nonnegative Francica20 , i.e., $\mathcal{E}_{c}(\rho)=\mathcal{E}(\rho)-\mathcal{E}(\Delta(\rho))\geq 0$. Then, the dephasing map is a free operation, $\Delta\in\mathcal{O}_{p}$. For the dephasing map, we get $T(\Delta(P)|\rho)=T(P|\rho)$, however we can get $T(\Delta(P)|\Delta(\rho))>T(P|\rho)$ for $d>2$. It is enough to consider $d=3$. From an explicit calculation, for $\rho=\frac{1}{2}(|\epsilon_{2}\rangle\langle\epsilon_{2}|+|\epsilon_{3}\rangle\langle\epsilon_{3}|+|\epsilon_{2}\rangle\langle\epsilon_{3}|+|\epsilon_{3}\rangle\langle\epsilon_{2}|)$, we find $P_{\rho}=|\epsilon_{1}\rangle\langle\epsilon_{1}|$, $\Delta(\rho)=\frac{1}{2}(|\epsilon_{2}\rangle\langle\epsilon_{2}|+|\epsilon_{3}\rangle\langle\epsilon_{3}|)$ and $P_{\Delta(\rho)}=\frac{1}{2}(|\epsilon_{1}\rangle\langle\epsilon_{1}|+|\epsilon_{2}\rangle\langle\epsilon_{2}|)$, thus we find that $T(\Delta(P)|\Delta(\rho))>T(P|\rho)$ for the passive states $P=\sum_{k}p_{k}|\epsilon_{k}\rangle\langle\epsilon_{k}|$ with populations $p_{k}\geq p_{k+1}$ satisfying the inequality
@@
We note that if $\Lambda$ is a unital map, then $E(P_{\Lambda(\rho)})\geq E(P_{\rho})$ (see, e.g., Ref. Francica22 ). Thus, if $E(\Lambda(\rho))\leq E(\rho)$, we get
@@
In particular, if $\Lambda(C)\subseteq C$, the condition $E(\Lambda(\gamma_{\beta}))\leq E(\gamma_{\beta})$ is equivalent to $T(\Lambda(\gamma_{\beta}))\leq T(\gamma_{\beta})$, i.e., $\eta_{\Lambda}(C)=1$. If $\Lambda$ is a unitary map defined by a unitary operator $U$, so that $\Lambda(\rho)=U\rho U^{\dagger}$, the condition (E,ii) is equivalent to $E(\Lambda(\rho))\leq E(\rho)$. For instance the map $\Lambda(\rho)=U_{\rho}\rho U_{\rho}^{\dagger}$ that extracts the ergotropy is free. Then, there is no a work locking phenomenon Horodecki13 ; Skrzypczyk14 ; Lostaglio15 , since all the ergotropy can be extracted with free unitary maps.
V Monotones
Let us introduce some monotones. In the following we go beyond the quantum relative entropy by focusing on the Tsallis divergence defined as
$S_{\alpha}(\rho||\sigma)=\frac{1}{\alpha-1}(\text{Tr}\left\{\rho^{\alpha}\sigma^{1-\alpha}\right\}-1)\,.$ (56)
Similarly to the quantum relative entropy, the ergotropy can be expressed as
$\mathcal{E}(\rho)=\min_{P\in\Pi}T_{\alpha}(P|\rho)S_{\alpha}(\rho||P)\,,$ (57)
which generalizes Eq. (30) that is achieved in the limit $\alpha\to 1$. We defined
$T_{\alpha}(P|\rho)=\frac{(1-\alpha)\mathcal{E}(\rho)}{\text{Tr}\left\{(P_{\rho}^{\alpha}-\rho^{\alpha})P^{1-\alpha}\right\}}\,,$ (58)
so that the minimum in Eq. (57) is obtained for $P=P_{\rho}$. Then we define the family of monotones
$M_{p,\alpha\nu}(\rho)=\min_{P\in\Pi}T^{\nu}_{\alpha}(P|\rho)S_{\alpha}(\rho||P)\,,$ (59)
where $\nu\in[0,1]$, since we have that $M_{p,\alpha\nu}(\rho)\geq M_{p,\alpha\nu}(\Lambda(\rho))$ for $\Lambda\in\mathcal{O}_{p}$.
Proof.
We aim to prove that $R_{\Lambda,\nu}(\rho)\leq 1$ for $\Lambda\in\mathcal{O}_{p}$, where we defined $R_{\Lambda,\nu}(\rho)=M_{p,\alpha\nu}(\Lambda(\rho))/M_{p,\alpha\nu}(\rho)$. We start to note that $R_{\Lambda,0}(\rho)\leq 1$ for all CPTP maps $\Lambda$ and $R_{\Lambda,1}(\rho)\leq 1$ for all $\Lambda\in\mathcal{O}_{p}$. In order to examine the case $\nu\in(0,1)$, we note that
$\partial_{\nu}M_{p,\alpha\nu}(\rho)=\frac{M_{p,\alpha\nu}(\rho)}{\nu}\ln\frac{M_{p,\alpha\nu}(\rho)}{M_{p,\alpha 0}(\rho)}\,,$ (60)
where $\partial_{\nu}M_{p,\alpha\nu}(\rho)|_{\nu=0}=M_{p,\alpha 0}(\rho)\ln(M_{p,\alpha 1}(\rho)/M_{p,\alpha 0}(\rho))$. The derivative of Eq. (60) can be calculated by considering the bound
$M_{p,\alpha\nu}(\rho)\geq M^{\frac{\nu}{x}}_{p,\alpha x}(\rho)M^{1-\frac{\nu}{x}}_{p,\alpha 0}(\rho)\,,$ (61)
where $x$ is a real number, which can be easily derived by noting that
$\min_{P\in\Pi}T^{\nu}_{\alpha}(P|\rho)S_{\alpha}(\rho||P)=\min_{P\in\Pi}\left(T^{x}_{\alpha}(P|\rho)S_{\alpha}(\rho||P)\right)^{\frac{\nu}{x}}S^{1-\frac{\nu}{x}}_{\alpha}(\rho||P)$
$\geq\min_{P\in\Pi}\left(T^{x}_{\alpha}(P|\rho)S_{\alpha}(\rho||P)\right)^{\frac{\nu}{x}}\min_{P\in\Pi}S^{1-\frac{\nu}{x}}_{\alpha}(\rho||P)\,.$ (62)
By considering $\epsilon\to 0$, for $x=\nu\pm\epsilon$ we get $M_{p,\alpha\nu}(\rho)\geq M_{p,\alpha\nu\pm\epsilon}(\rho)(M_{p,\alpha\nu\pm\epsilon}(\rho)/M_{p,\alpha 0}(\rho))^{\mp\frac{\epsilon}{\nu}}=M_{p,\alpha\nu\pm\epsilon}(\rho)\mp\epsilon M_{p,\alpha\nu}(\rho)\ln(M_{p,\alpha\nu}(\rho)/M_{p,\alpha 0}(\rho))/\nu+\mathcal{O}(\epsilon^{2})$, from which it follows Eq. (60). Thus, from Eq. (60) we get
$\partial_{\nu}R_{\Lambda,\nu}(\rho)=\frac{R_{\Lambda,\nu}(\rho)}{\nu}\ln\frac{R_{\Lambda,\nu}(\rho)}{R_{\Lambda,0}(\rho)}\,,$ (63)
so that $\partial_{\nu}R_{\Lambda,\nu}(\rho)\geq 0$ when $R_{\Lambda,\nu}(\rho)\geq R_{\Lambda,0}(\rho)$, whereas $\partial_{\nu}R_{\Lambda,\nu}(\rho)\leq 0$ when $R_{\Lambda,\nu}(\rho)\leq R_{\Lambda,0}(\rho)$. Let us defines a general partition of $[0,1]$ such that $\partial_{\nu}R_{\Lambda,\nu}(\rho)\leq 0$ for $\nu\in[\nu_{2i-1},\nu_{2i}]$ and $\partial_{\nu}R_{\Lambda,\nu}(\rho)\geq 0$ for $\nu\in[\nu_{2i},\nu_{2i+1}]$, where $\nu_{i}$ are such that $0\leq\nu_{i}\leq\nu_{i+1}\leq 1$. For all $\Lambda\in\mathcal{O}_{p}$, for $\nu\in[\nu_{2i-1},\nu_{2i}]$ we get $R_{\Lambda,\nu}(\rho)\leq R_{\Lambda,0}(\rho)\leq 1$, whereas for $\nu\in[\nu_{2i},\nu_{2i+1}]$ we get $R_{\Lambda,\nu}(\rho)\leq R_{\Lambda,\nu_{2i+1}}(\rho)\leq 1$, where $R_{\Lambda,\nu_{2i+1}}(\rho)\leq 1$ since $\nu_{2i+1}\in[\nu_{2i+1},\nu_{2i+2}]$ or $\nu_{2i+1}=1$ and $R_{\Lambda,1}(\rho)\leq 1$. For instance, let us consider $\nu_{1}=0$ and $\nu_{5}=1$, we get $R_{\Lambda,\nu}(\rho)\leq 1$ for $\nu\in[\nu_{1},\nu_{2}]\cup[\nu_{3},\nu_{4}]$, $R_{\Lambda,\nu}(\rho)\leq R_{\Lambda,\nu_{5}}(\rho)=R_{\Lambda,1}(\rho)\leq 1$ for $\nu\in[\nu_{4},\nu_{5}]$ and $R_{\Lambda,\nu}(\rho)\leq R_{\Lambda,\nu_{3}}(\rho)\leq 1$ for $\nu\in[\nu_{2},\nu_{3}]$ since $\nu_{3}\in[\nu_{3},\nu_{4}]$ and thus $R_{\Lambda,\nu_{3}}(\rho)\leq 1$, then $R_{\Lambda,\nu}(\rho)\leq 1$ for all $\nu\in[0,1]$ if $\Lambda\in\mathcal{O}_{p}$. ∎
Similarly, concerning the free energy we get
$F(\rho)=\min_{\gamma\in C}T_{cp,\alpha}(\gamma|\rho)S_{\alpha}(\rho||\gamma)\,,$ (64)
where we defined
$T_{cp,\alpha}(\gamma|\rho)=\frac{(1-\alpha)F(\rho)}{\text{Tr}\left\{(\gamma_{\rho}^{\alpha}-\rho^{\alpha})\gamma^{1-\alpha}\right\}}\,.$ (65)
Then we achieve the family of monotones
$M_{cp,\alpha\nu}(\rho)=\min_{\gamma\in C}T^{\nu}_{cp,\alpha}(\gamma|\rho)S_{\alpha}(\rho||\gamma)\,,$ (66)
where $\nu\in[0,1]$, such that $M_{cp,\alpha\nu}(\rho)\geq M_{cp,\alpha\nu}(\Lambda(\rho))$ for $\Lambda\in\mathcal{O}_{cp}$. In the end, we note that (v) the additivity property is never satisfied since if $\Lambda$ is a free operation of these resource theories, the operation $\Lambda\otimes\text{id}$ is not free on the Hilbert space $\mathcal{H}\otimes\mathcal{H}^{\prime}$, where id is the identity map on the Hilbert space $\mathcal{H}^{\prime}$, i.e., the free operations are not completely free operations. Basically, the tensor product of two free states globally is not always a free state. Due to this tensor structure, general results derived for resource theories (see, e.g., the results in Ref. Liu19 ) cannot be applied. As usual for these monotones, if an operation satisfies (iii) the strong monotonicity and (iv) the convexity properties then it is expected to also satisfy (ii) the monotonicity property.
VI Conclusions
@@
Quantum resource theories are typically useful to identify processes for the utilization of the resource for a given task. Here, we took in exam the work extractable as a possible resource. Work extraction in thermally isolated quantum systems can be performed through unitary cycles. The maximum work extractable is the ergotropy, which is upper bounded by the free energy. When we have many copies of the system available, the density of maximum work extractable tends to be the free energy. Thus, we defined resource theories where the free energy and the ergotropy are resources and thus monotones. Completely passive states and passive states are the free states of the resource theories. In particular, we discussed how, similarly to the free energy, the ergotropy is a distance-based monotone. Thus, with the aim to give a unified picture, we expressed both the free energy and the ergotropy in terms of a unique metric, with a non-contractive distance (in broad sense). This distance is defined in terms of the quantum relative entropy and a sort of nonequilibrium temperature, which plays a certain role to characterize the free operations. In detail, any map for which the distance is contractive is a free operation for the corresponding theory. However, there are also free operations for which the distance is not contractive, finding that all the free operations can increase the temperature of the free states at most by a certain factor not smaller than 1. To summarize, thanks to the introduced distance we are able to identify the free operations based on how they act on the free states.
@@
In conclusion, we hope that our results can be useful to understand how the work extractable can be viewed as a resource and can also find applications for the manipulation of the states of work storing centers, e.g., quantum batteries. Among the possible applications, we note that the charging of quantum batteries, which increases the work extractable of the batteries, can be performed only through some not free operation. Furthermore, we believe that our findings can also play some role in understanding what is the temperature of a nonequilibrium state. In particular, the definition of the nonequilibrium temperature is a challenging open problem of quantum thermodynamics, e.g., a nonequilibrium temperature was recently introduced in Ref. Alipour21 , which is different from the functional introduced by us.